The perimeter of a rectangular field is 294 yards. If the length of the field sis 93 yards, what is the width? I came up with 54 yards, but could someone check over my work, thanks:)
Answer by checkley77(12844)

2*93+2W=294
186+2W=294
2W=294-186
2W=108
W=108/2
W=54 IS THE WIDTH.